1. A model rocket has a mass of 0.435 kg. It is mounted vertically and launches directly up with an thrust force of 13.1 N. What is the acceleration of the rocket after liftoff?
1
Expert's answer
2021-04-26T17:01:42-0400
"a=\frac{T-W}{m}\\a=\frac{T}{m}-g\\a=\frac{13.1}{0.435}-9.8=20.3\frac{m}{s^2}"